Solve by the addition method. (If there is no solution, enter NO SOLUTION. Use the parameters x and y as necessary.)
x + 5y = 12
2x + 7y = 15

Answer by Edwin McCravy(20056)

Instead of doing it for you, I'll do one exactly like it step-by-step.
Use it as a model for doing your problem.

 x + 7y = 24
4x + 9y = 20

Multiply the first equation through by -4 to make the x become -4x
so it will cancel out when adding it to 4x in the second equation.

-4x - 28y = -96
 4x +  9y =  20

Then we add them term by term:

-4x - 28y = -96
 4x +  9y =  20
---------------
     -19y = -76
        y = (-76)/(-19)
        y = 4

Substitute 4 for y in

   x + 7y = 24
 x + 7(4) = 24
   x - 28 = 24
        x = -4

Now do yours exactly the same way step by step.

Edwin